Help please

Workshop 5 Problems Student Q Search Sheet E 2 A OBOU Home Insert Page Layout Formulas Data Review View , Calibri (Body) -
Answer #1
  • Requirement, as asked

Assets

Stockholder's equity

Income Statement

Cash

Land

Common Stock

PIC in excess

Revenues

Expenses

Net Income

a

$20,000

$10,000

$10,000

b

$50,000

$20,000

$30,000

--Working

Assets

Stockholder's equity

Income Statement

Cash

Land

Common Stock

PIC in excess

Revenues

Expenses

Net Income

a

20000

=1000*10

=20000-10000

b

=2000*25

=2000*10

=50000-20000